# Look-ahead Search on Top of Policy Networks in Imperfect Information Games

### 23 Dec 2023 v1

@@ 4.3. Value Function
Figure 2. Modified subgame with multi-valued states, that contains 3 leaf public states. Right part shows detail of public state $s_{0}^{1}$ that contains 4 histories and 2 information sets for each player. The multi-valued states technique modifies this public state by giving player 2 a choice between two strategies in each history against blueprint policy of player 1.
@@
Multi-valued states Brown et al. (2018) are used as a modification of the original game, where after the depth limit, the opponent has one more decision point, in which it chooses its strategy for the remainder of the game against some fixed blueprint policy. This modified game is a standard game and can be solved by any game solving algorithm unaware of value functions. Authors of Moravčík et al. (2017) trained value function for Poker, which served as a generalization of the value function from perfect information games. This value function returns counterfactual values based on ranges at the depth limit. These were then specially adapted for Counterfactual Regret Minimization.
@@
These two approaches to the depth-limited search may not seem connected, but the authors of Kovařík et al. (2023) show this connection by explaining that the multi-valued states could be viewed as a value function, which is not trained for all possible strategies in the game, but only for some subset. However, their explanation does not explicitly clarify how this connection could be leveraged.
@@
Let us assume a depth-limited subgame, as in Figure 2, with three leaf public states $s_{0}^{1},s_{0}^{2},s_{0}^{3}$. Further examining public state $s_{0}^{1}$ in Figure 1(b), there are four possible states in this public state, which are separated into different infosets. First player has infoset $s_{1}^{1}$, that contains histories $h_{1},h_{3}$ and $s_{1}^{2}$ containing historis $h_{2},h_{4}$. Similarly second player has infosets $s_{2}^{1},s_{2}^{2}$, that contains histories $h_{1},h_{2}$ and $h_{3},h_{4}$ respectively. The value function would return counterfactual values for each infoset, namely $v^{\pi}_{1}(s_{1}^{1}),v^{\pi}_{1}(s_{1}^{2}),v^{\pi}_{2}(s_{2}^{1}),v^{\pi}_{2}(s_{2}^{2})$.
@@
In the same example, using multi-valued states as proposed in Brown et al. (2018) gives the opponent, which is player 2, another choice to pick between two different strategies in each infoset, as shown in Figure 1(b). The counterfactual values are in this case computed as $v^{\pi}_{i}(s_{i})=\frac{\sum_{h\in s_{i}}P_{-i}(h)\sum_{a\in\mathcal{A}(s_{2}(h))}\pi_{2}(s_{2}(h),a)\mathcal{R}_{i}(h,a)}{\sum_{h\in s_{i}}P_{-i}(h)}$ for each player $i$ and it’s infosets $s_{i}\in\mathcal{S}_{i}(s_{0})$. Reward in this case corresponds to the value taken from the critic $\mathcal{R}_{i}(h,a)=\boldsymbol{u}_{i}(h)[f^{T}]$, where each action $a$ corresponds to a single policy transformation $f^{T}$.
@@
Performing Counterfactual Regret Minimization (CFR) requires storing multiple statistics in each infoset, like regret and average strategy. These statistics must be stored even for these artificial infosets created by the multi-valued states. In such a scenario, the multi-valued states could be viewed as a stateful value function where these statistics determine the value function state. Each CFR iteration, this value function updates its state and returns counterfactual values based on its current state and input ranges $r$. However, let us assume that instead of performing regret matching and storing these necessary statistics, the opponent would always pick a best response based on input ranges. In such a case, we could use multi-valued states as a usual value function with just additional computation step, without keeping the state inside the value function. Let us rephrase the value function in terms of using a history network with policy transformations for player $i$.
@@
With the constructed gadget game and a corresponding value function, any previously developed search algorithm for imperfect information games may conduct the subsequent search. During our experiments, we used CFR+ Tammelin et al. (2015) exclusively. When using CFR+, a subclass of Growing tree-CFR (GT-CFR) introduced in Schmid et al. (2023), the bounds already established by the GT-CFR authors in Theorem 3 also apply to our approach. However, the quality of this value function $\xi$, which is the distance between values returned by the optimal value function and the trained value function, is affected by additional errors, even if the value function is trained perfectly. The first of these errors is that the blueprint strategy of resolving player $i$ is not Nash equilibrium but just some approximation. Second is that the opponent’s strategies used for multi-valued states may not be best responses to any strategy. As stated by authors of Brown et al. (2018) in Proposition 1, if the blueprint strategy is Nash equilibrium, the opponent strategies are all pure actions, then the resulting strategy is also part of a Nash equilibrium.
@@
The multi-valued states could be expanded by giving both players a choice to pick from multiple strategies for the rest of the game Brown and Sandholm (2019); Kovařík et al. (2023). This would increase the number of values that must be trained quadratically because each pair of strategies between players yields different values. In Figure 2, the opponent can choose which of the two strategies it will follow for the rest of the game against the blueprint. If we would give both players a choice to pick between two strategies to follow, then there would be four joint actions as a combination between these two strategies. Multi-agent Search with Policy Transformation Value Network is compatible with this approach, and the only required change is to compute a Nash equilibrium of the underlying normal-form game at the depth limit during search instead of a best response.
@@
We have already shown how the history critic trained for multi-valued states can be used as a value function. However, we could go one step further and train the usual value function out of this history critic. This would completely avoid the need to call the critic for each history and further compute a best response or Nash equilibrium when using multi-valued states for single or both players. The value function may be trained before the gameplay phase by randomly sampling possible ranges, computing the counterfactual values, and then training the value function on these counterfactual values. Such trained value function can then be used in already developed algorithms Moravčík et al. (2017); Schmid et al. (2023) without any changes. The quality of such a value function depends on the chosen transformations when training multi-valued states. In our work, we have always used the history critic and have done computation of a best response during the search.
5. Experiments
We have conducted three main experiments to validate our approach’s effectiveness. In the first experiment, we compare head-to-head play in larger games to validate that the inclusion of search improves the strategy from the policy network. Head-to-head play may not be a good indicator of strategy quality in imperfect information games, so we further conducted experiments to show that using search decreases overall exploitability even in smaller games. The exploitability as a metric shows how well a best-responding opponent can play against the policy. Strategies with small exploitability are, therefore, robust against any opponent. The third experiment shows how computing counterfactual values from the history critic in gadget games affects the quality of the solution. Appendix B provides a detailed description of the game rules. All Goofspiel and Battleships experiments have used the same 22 transformations that are described in Appendix C, along with other hyperparameters. All of our experiments have used the framework OpenSpiel Lanctot et al. (2019). The depth-limited search was conducted on a single core of the CPU Intel Xeon Scalable Gold 6146, operating at a frequency of 3.2 GHz, while the network training additionally used a single GPU Tesla V100.
5.1. Head-to-head play in large games
@@
Our first experiment shows the scalability and effectiveness of SePoT on large games. We separately trained RNaD and RNaD with SePoT, each three times, with the same hyperparameter setting. These were then used in head-to-head play against each other, where RNaD directly followed the policy network and SePoT applied search in those subgames that contained less than 25000 or 100000 unique histories. The evaluation encompassed these games: Goofspiel with eight cards and descending order of point cards, which contains approximately $2\cdot 10^{9}$ unique histories; Goofspiel with eight cards and randomized order of point cards, having $9\cdot 10^{13}$ unique histories; Goofspiel with thirteen cards and descending order of point cards, containing approximately $5\cdot 10^{19}$ unique histories; Goofspiel with thirteen cards and randomized order of point cards, having roughly $3\cdot 10^{29}$ unique histories; Battleships on a square grid of size 3 with 2 ships of sizes 2 and 2 containing roughly $10^{14}$ unique histories and Battleships on square grid of size 5 with 3 ships of sizes 3,2 and 2 that have roughly $10^{57}$ unique histories.
@@
For each game, we have trained 3 RNaD policy networks and 3 different RNaD policy networks along with SePoT’s history critic. Each network was trained for 3 days. We have run over one million game simulations of RNaD against SePoT for each game. In half of them, the SePoT was a first player; in half, it was a second player. The average rewards multiplied by 100 from these simulations are presented in Table 1 along with a 95% confidence interval. Employing search outperformed the RNaDs policy network with statistical significance in all the games tested.
Increasing the size of subgames that we use search in slightly improves the performance of SePoT in almost all cases, which shows that incorporating additional search leads to better results. The only exception is a Goofspiel 13 with descending point card ordering. This is most likely because there are only slightly more subgames in which the search is used with this increased limit, and in those, the RNaDs policy network performs well. The usage of search improves the performance in larger Battleships much more than in any other game. This is because, in Goofspiel, most of the subgames are quite large, so the search is used sparingly. However, Battleships have a lot of smaller subgames where the search is used. Moreover, the Battleships is a much larger game than the Goofspiel, and similar subgames in Battleships may have vastly different optimal policies. This suggests that training RNaD in Battleships is harder than in Battleships. On the other hand, the smaller Battleships are small enough that the policy network could be sufficiently trained, and the performance of search there is almost on par with the policy network.
@@
25000 $1.42\pm 0.01$ $3.71\pm 0.03$ $6.31\pm 0.01$ $1.17\pm 0.01$ $0.44\pm 0.06$ $34.58\pm 0.16$
100000 $1.66\pm 0.02$ $4.47\pm 0.11$ $6.28\pm 0.08$ $1.47\pm 0.16$ $2.83\pm 0.06$ $36.54\pm 0.2$
Table 1. Average rewards from head-to-head play of RNaD against RNaD with SePoT. The average rewards are multiplied by 100.
5.2. Exploitability in smaller games
In the second experiment, we aim to evaluate the effect of depth-limited search with multi-valued states on the exploitability of the resulting policy compared to the policy network. In the experiments, we compare the exploitability of the policy network from RNaD and the SePoT when using different look-ahead lengths. Firstly, we compared the exploitability in Leduc hold’em, where we split the game on the part before dealing a public card and after. We used constant transformations that transformed any policy into each pure action. We trained ten networks with different random seeds while computing exploitability every 2,000 iterations up to iteration 1,000,000. The results in Figure 2(a) show the mean and 95% confidence interval of the exploitability throughout these runs with increasing iterations. We see that SePoT cannot exceed the performance of the RNaD in Leduc hold’em, but during the training the SePoT seems to be more stable.
We conducted the same experiment in Goofspiel with five cards and randomized point card ordering for 100,000 RNaD iterations. This game consists of 2.4 million unique histories. We used different depth limits for search, where DL=$n$ means that both players play $n$ actions before the value function is called from the root of the current subgame. The plots in Figure 2(b) show how the exploitability changes with additional iterations. Using the depth limit 2 or 3 yields similar results that both outperform depth limit 1 and the policy network. This experiment shows that with additional training, the search still improves over the network policy. Furthermore, it confirms the intuition that having a longer look-ahead increases the quality of the search solution.
Figure 3. Exploitability based on the RNaD training comparing policy network from RNaD and the search with SePoT.
@@ 5.2. Exploitability in smaller games
Figure 3. Exploitability based on the RNaD training comparing policy network from RNaD and the search with SePoT.
5.3. Counterfactual values in gadget games
Our final experiment aims to show that using the counterfactual values computed from history critic does not significantly worsen the solution compared to using them from previous resolves. We trained ten critic networks with the same hyperparameter setting for Leduc hold’em and Goofspiel with five cards and randomized point card ordering. Then, we solved each subgame separately, either by computing the counterfactual values for the gadget game from the history critic or reusing them from previous searches. In Leduc hold’em, we split the game before dealing a public card and after. In Goofspiel, each subgame had a depth limit of 1, so the value function is called after a single action by each player from the root of the subgame. The action in the gadget game is not counted towards the depth limit. We then combine these subgame policies into a policy for a whole game and compute the exploitability.
The results are shown in Figure 4, where we show the mean and the 95% confidence interval for both approaches. Using the counterfactual values from previous searches significantly outperforms using the history critic. On the other hand, using the counterfactual values in Goofspiel only slightly improves the performance of the search. In large games, where we cannot do all the previous searches, using the history critic is the only possibility to perform a search. This experiment shows that using the history critic may harm the final solution in some games. With this insight, the remainder of our experiments were conducted with the computation of the counterfactual values from the critic network.
Figure 4. Exploitability based on the training iterations, when using the counterfactual values from previous searches or computing them from history critic.
(a) Leduc hold’em
(b) Goofspiel
Goofspiel and Leduc Exploitability
Figure 4. Exploitability based on the training iterations, when using the counterfactual values from previous searches or computing them from history critic.
6. Related work
The model-free policy-gradient techniques, which serve as a main building block of our algorithm, are applicable to huge imperfect information games, with the only limitation being the size of the player’s private knowledge that serves as an input to the neural network. However, safe search is a much more limited technique since it requires a model of the game to construct the subgame. Moreover, safe search is much more computationally demanding, since it requires storing the subgame in a memory.
Some techniques to overcome this issue include either abstracting the game Brown and Sandholm (2018) or revealing part of the private information Zhang and Sandholm (2021). These methods decrease the size of the subgame at the cost of the final solution. Since previous algorithms use search in both train and test time Brown et al. (2020); Brown and Sandholm (2018); Moravčík et al. (2017); Schmid et al. (2023), it becomes necessary to use these techniques to downsize some subgames in huge games, even to be able to train the value functions and apply the search algorithms.
@@ 6. Related work
In our work, we use a policy-gradient training algorithm that trains at arbitrarily sized games. We can choose whether the search should be used at each point independently during gameplay. This enables our search in larger games than previously possible without using these subgame reduction techniques. However, SePoT is flexible enough to be used in unison with these subgame reduction techniques to increase the number of subgames that the search may be applied in even further.
@@
Imagine a game of Battleships as an example. After placing the ships, there are more than $10^{20}$ possible states in a public state. Safe search in such a subgame is infeasible. However, after several more rounds, the information is steadily being revealed to both players, and eventually, the size of the public state gets small enough. Our proposed algorithm can perform search in such a situation, which was previously not possible.
@@
Our method is heavily based on the Brown et al. (2018), with the notable differences being a more principled way to train the blueprint policy with policy-gradient algorithms, more thoroughly connecting the multi-valued states with value functions. We show how the multi-valued states may be used to train the value function or how to use them as a value function. Furthermore, we discuss how the multi-valued states could be expanded to the setting where both players can choose their strategy until the end of the game. Lastly, we present a way to further leverage an already trained critic network to construct a gadget game anywhere in the game without the need to perform search in previous parts.
We have exclusively used RNaD Perolat et al. (2022) as a policy-gradient training algorithm. SePoT could be trained alongside any different multi-agent reinforcement learning algorithms for imperfect-information games that approximate Nash Equilibrium from sampled trajectories. Since SePoT trains on the same trajectories, it uses them more effectively. SePoT could be used even with an already trained policy, but the training then requires sampling additional trajectories, which is not always possible.
Since the search is more computationally costly than following the policy from the neural network, with SePoT it is possible to argue when to perform search to improve the gameplay performance only in important parts of the game according to some heuristic.
@@
Model-free deep reinforcement learning methods based on regret minimization Steinberger et al. (2020); Gruslys et al. (2020); McAleer et al. (2023) use search in the gameplay and still can be scaled to very large games similarly to SePoT. The main limitation of these methods is the need to compute the average strategy, which either requires storing all the trajectories used in learning or approximating this average, which is prone to error. In games like Battleships, the training may consist of hundreds of millions of trajectories to train properly. Storing all of these trajectories is memory-demanding, and computing the average out of them is also computation-demanding. SePoT uses a policy-gradient algorithm that approximates the optimal strategy in immediate strategies, so it does not require computing the average strategy. For search in the gameplay phase, all the necessary statistics are stored within the trained policy network and history critic.
Algorithms for imperfect information games outside of two-player zero-sum scope like Pluribus for multiplayer Poker Brown and Sandholm (2019), Sparta for cooperative game Hanabi Lerer et al. (2020) or Cicero for the multiplayer adversarial game with cooperative elements Diplomacy (FAIR)† et al. (2022) all used search to achieve strong performances. These advancements underscore the significance of search as an integral component in gameplaying algorithms.
7. Conclusion

### 06 Feb 2024 v2

@@ 4.3 Value Function
Figure 2: Subgame with 3 leaf public states. Right part shows the detail of public state $s_{0}^{1}$ with 4 histories and 2 information sets for each player. The multi-valued states technique modifies this public state by giving player 2 a choice between two strategies in each history against blueprint policy of player 1.
@@
Multi-valued states (Brown et al., 2018) are a modification of the original game, where after the depth limit, the opponent has one more decision point, in which it chooses its strategy for the remainder of the game against some fixed blueprint policy. This modified game is a standard game and can be solved by any game solving algorithm unaware of value functions. Authors of (Moravčík et al., 2017) trained value function for Poker, which served as a generalization of the value function from perfect information games. This value function returns counterfactual values based on ranges at the depth limit. These were then specially adapted for Counterfactual Regret Minimization (CFR) (Zinkevich et al., 2007).
@@
These two approaches to the depth-limited search may not seem connected, but the authors of (Kovařík et al., 2023) show this connection by explaining that the multi-valued states could be viewed as a value function, which is not trained for all possible strategies in the game, but only for some subset. However, their explanation does not explicitly clarify how this connection could be leveraged.
@@
Let us assume a depth-limited subgame, as in Figure 2, with three leaf public states $s_{0}^{1},s_{0}^{2},s_{0}^{3}$. Further examining public state $s_{0}^{1}$ in Figure 1(b), there are four possible states in this public state, which are separated into different infosets. First player has infoset $s_{1}^{1}$, that contains histories $h_{1},h_{3}$ and $s_{1}^{2}$ containing histories $h_{2},h_{4}$. Similarly second player has infosets $s_{2}^{1},s_{2}^{2}$, that contains histories $h_{1},h_{2}$ and $h_{3},h_{4}$ respectively. The value function would return counterfactual values for each infoset, namely $v^{\pi}_{1}(s_{1}^{1}),v^{\pi}_{1}(s_{1}^{2}),v^{\pi}_{2}(s_{2}^{1}),v^{\pi}_{2}(s_{2}^{2})$.
@@
In the same example, using multi-valued states as proposed in (Brown et al., 2018) gives the opponent, player 2, another choice to pick between two strategies in each infoset, as shown in Figure 1(b).
@@
Using CFR requires storing multiple statistics in each infoset, like regret and average strategy. These statistics must be stored even for these artificial infosets created by the multi-valued states. In such a scenario, the multi-valued states could be viewed as a stateful value function where these statistics determine the value function state. Each CFR iteration, this value function updates its state and returns counterfactual values based on the current state and input ranges $r$. Let us assume that instead of performing regret matching and storing these necessary statistics, the opponent always picks a best response based on input ranges. In such a case, we can use multi-valued states as a usual value function with just additional computation step, without keeping the state inside the value function. Let us rephrase the value function in terms of using a history network with policy transformations for player $i$.
@@
This presents a new way to train value functions without using search in the training time at all by first training multi-valued states and then using them with different possible ranges to generate counterfactual values for training the value function.
@@
With the constructed gadget game and a corresponding value function, any previously developed search algorithm for imperfect information games may conduct the subsequent search. During our experiments, we used CFR+ (Tammelin et al., 2015) exclusively. When using CFR+, a subclass of Growing tree-CFR (GT-CFR) introduced in (Schmid et al., 2023), the bounds already established by the GT-CFR authors in Theorem 3 also apply to our approach. However, the quality of this value function $\xi$, which is the distance between values returned by the optimal value function and the trained value function, is affected by additional errors, even if the value function is trained perfectly. The first of these errors is that the blueprint strategy of resolving player $i$ is not Nash equilibrium but just some approximation. Second is that the opponent’s strategies used for multi-valued states may not be best responses to any strategy.
@@
The multi-valued states could be expanded by giving both players a choice to pick from multiple strategies for the rest of the game (Brown and Sandholm, 2019; Kovařík et al., 2023). This would increase the number of values that must be trained quadratically because each pair of strategies between players yields different values. SePoT is compatible with this approach, and the only required change is to compute a Nash equilibrium of the underlying normal-form game at the depth limit during search instead of a best response.
@@
We have already shown how the history critic trained for multi-valued states can be used as a value function. However, we could go one step further and train the usual value function out of this history critic. The value function may be trained before the gameplay phase by randomly sampling possible ranges, computing the counterfactual values from multi-valued states, and then training the value function on these counterfactual values. Such trained value function can then be used in already developed algorithms (Moravčík et al., 2017; Schmid et al., 2023) without any changes.
We provide pseudocode for training and gameplay in Appendix E
5 Experiments
We have conducted two experiments to validate SePoT’s performance. Our experiments have used the framework OpenSpiel (Lanctot et al., 2019). The depth-limited search was conducted with CPU Intel Xeon Scalable Gold 6146, operating at a frequency of 3.2 GHz, while the network training additionally used a single GPU Tesla V100. The hyperparameter setting is described in Appendix D. Appendix B provides a detailed description of the game rules.
5.1 Head-to-head play in large games
@@
The first experiment shows the effectiveness of SePoT in large games. We trained RNaD and RNaD with SePoT separately, each three times, with the same hyperparameter setting. These were then used in head-to-head play against each other, where RNaD directly followed the policy network and SePoT applied search in those subgames that contained less than 25000 or 100000 unique histories. The evaluation encompassed these games: Goofspiel with eight cards and randomized point order, and with thirteen cards both with descending and randomized point order and Battleships on square grids of size 3x3, 5x5, and 7x7 with ships of size (2, 2), (3, 2, 2), (4, 3, 3, 2) respectively.
@@
For each game, we have trained 3 RNaD policy networks and 3 different RNaD policy networks along with SePoT’s history critic. Each network was trained for 3 days. We have run over one hundred thousand game simulations of RNaD against SePoT for each game. In half of them, the SePoT was a first player; in half, it was a second player. The average rewards multiplied by 100 from these simulations are presented in Table 1 along with a 95% confidence interval. Employing search outperformed the RNaDs policy network with statistical significance in all the games tested.
The usage of search improves the performance in Battleships more than in Goofspiel. This is because, in Goofspiel, most of the subgames are quite large, so the search is used sparingly. However, Battleships have a lot of smaller subgames where the search is used.
@@
Battleships
7x7
25000 $2.70\pm 0.05$ $3.44\pm 0.05$ $1.12\pm 0.05$ $6.2\pm 0.06$ $81.9\pm 0.3$ $24.1\pm 0.7$
100000 $3.10\pm 0.19$ $3.51\pm 0.1$ $1.43\pm 0.3$ $6.5\pm 0.07$ $83.5\pm 0.5$ $42.1\pm 0.9$
Table 1: Average rewards from head-to-head play of RNaD against RNaD with SePoT. The average rewards are multiplied by 100.
5.2 Exploitability in smaller games
In the second experiment, we aim to evaluate the effect of depth-limited search with multi-valued states on the quality of the policy compared to the policy network. In the experiments, we compare the exploitability of the policy network from RNaD and the SePoT. We use Leduc hold’em and Goofspiel with 5 cards and randomized order of point cards for evaluation. We trained ten networks with different random seeds while computing exploitability every 500 iterations up to iteration 75,000. The results in Figure 3 show the mean and 95% confidence interval of the exploitability throughout these runs. In Leduc hold’em, we separately solve part before dealing a public card. In Goofspiel, we use different depth limits for search, where DL=$n$ means that both players play $n$ actions before the value function is called from the root of the current subgame.
The plots in Figure 2(b) show how the exploitability changes with additional iterations. Using the depth limit 2 or 3 yields similar results, showing that both outperform depth limit 1 and the policy network. This experiment shows that with additional training, the search still improves over the network policy. Furthermore, it confirms the intuition that having a longer look-ahead increases the quality of the search solution.
@@ 5.2 Exploitability in smaller games
Figure 3: Exploitability of policy network RNaD and the search with SePoT based on RNaD training iterations.
6 Related work
@@
Previous search algorithms require using safe search both in train and test time (Brown et al., 2020; Brown and Sandholm, 2018; Moravčík et al., 2017; Schmid et al., 2023). Using those algorithms in larger games requires some subgame reduction techniques like abstracting the game (Brown and Sandholm, 2018) or revealing part of the private information (Zhang and Sandholm, 2021). SePoT uses a policy-gradient training algorithm that trains in arbitrarily sized games. Furthermore, SePoT chooses whether the search should be used at each point independently during gameplay. This enables safe search in larger games than previously possible without using any subgame reduction techniques. However, SePoT is flexible enough to be used with these subgame reduction techniques to increase the number of subgames that the search may be applied in even further.
@@
Imagine a game of Battleships as an example. After placing the ships, there are more than $10^{20}$ possible states in a public state, which makes training of search algorithms infeasible. After several more rounds, the information is steadily being revealed to both players, and eventually, the size of the public state becomes small enough. SePoT can perform search in such a game, which was previously not possible.
@@
Our method is heavily based on the (Brown et al., 2018), with the notable differences being a more principled way to train the blueprint policy with policy-gradient algorithms and choosing the policy transformations. Moreover, we show how to use multi-valued states as a value function practically. We present a way to leverage an already trained critic network to construct a gadget game anywhere in the game without the need to perform a search in previous parts.
@@
Model-free deep reinforcement learning methods based on regret minimization (Steinberger et al., 2020; Gruslys et al., 2020; McAleer et al., 2023) use search in the gameplay and still can be scaled to very large games. The main limitation of these methods is the need to compute the average strategy, which either requires storing all the trajectories used in learning or approximating this average, which is prone to error. In games like Battleships, the training may consist of billions of trajectories to train properly. Storing all of these trajectories and computing averages is both memory and computationally demanding. SePoT uses a policy-gradient algorithm that approximates the optimal strategy in immediate strategies, so it does not compute the average strategy.
7 Conclusion